An object moves in a straight line starting from rest. There are two stage in The journey
1.it gain speed uniformly 2.0 seconds and attains a speed of 8.0ms-¹
2. It continues at this pace for a further 1.5 second. Sketch the graph of speed against time
Find (1)the acceleration in stage (a)
(2)the acceleration in stage (b)
(3)the total distance moved during stages (a)and (b)
1
Expert's answer
2020-10-27T11:16:33-0400

1)


"a=\frac{v}{t}=\frac{8}{2}=4\frac{m}{s^2}"

2)


"a=0\frac{m}{s^2}"

3)


"s=0.5at^2+vt'=0.5(4)(2)^2+(8)(1.5)=20\ m"